### v1

# PathSeeker: Exploring LLM Security Vulnerabilities with a Reinforcement Learning-Based Jailbreak Approach

Abstract.
Warning:This paper contains harmful LLM responses that may make you uncomfortable.
In recent years, Large Language Models (LLMs) have gained widespread use, accompanied by increasing concerns over their security. Traditional jailbreak attacks rely on internal model details or have limitations when exploring the unsafe behavior of the victim model, limiting their generalizability. In this paper, we introduce PathSeeker, a novel black-box jailbreak method inspired by the concept of escaping a security maze. This work is inspired by the game of rats escaping a maze. We think that each LLM has its unique “security maze”, and attackers attempt to find the exit learning from the received feedback and their accumulated experience to compromise the target LLM’s security defences. Our approach leverages multi-agent reinforcement learning, where smaller models collaborate to guide the main LLM in performing mutation operations to achieve the attack objectives. By progressively modifying inputs based on the model’s feedback, our system induces richer, harmful responses. During our manual attempts to perform jailbreak attacks, we found that the vocabulary of the response of the target model gradually became richer and eventually produced harmful responses. Based on the observation, we also introduce a reward mechanism that exploits the expansion of vocabulary richness in LLM responses to weaken security constraints. Our method outperforms five state-of-the-art attack techniques when tested across 13 commercial and open-source LLMs, achieving high attack success rates, especially in strongly aligned commercial models like GPT-4o-mini, Claude-3.5, and GLM-4-air with strong safety alignment. This study aims to improve the understanding of LLM security vulnerabilities and we hope that this sturdy can contribute to the development of more robust defenses.
copyright: none
1. Introduction
Large Language Models (LLMs) demonstrate the potential for general artificial intelligence (Bubeck et al., 2023). However, LLMs also face security threats (Yao et al., 2024b). Currently, the primary approach to addressing LLM security threats is through safety alignment techniques (Bai et al., 2022, Zhao et al., 2024, Song et al., 2024), ensuring that the output of LLMs aligns with human ethics and values. To test the security of LLMs, various jailbreak techniques (Jin et al., 2024a) have been proposed, such as GCG (Zou et al., 2023), CodeChameleon (Lv et al., 2024), GPTFuzzer (Yu et al., 2023), RL-JACK (Chen et al., 2024) and RLTA (Wang et al., 2024). These techniques can be categorized into white-box attacks, which rely on internal model information, and black-box attacks, which do not. Black-box attacks, in particular, do not require access to the model’s internal details, making them applicable to a wide range of models. Most black-box attack algorithms depend on LLMs analyzing the output of the target model to improve the next attack. GPTFuzzer (Yu et al., 2023), RL-JACK (Chen et al., 2024) and RLTA (Wang et al., 2024) are most close to our work while we are definitely different, and our solution is more general. GPTFuzzer uses a fine-tuned small model in the attack loop but does not leverage the experience gained from each attack attempt. RL-JACK and RLTA aim to improve the effectiveness of black-box attacks by incorporating intermediate feedback during the attack process. However, both approaches rely on harmful reference answers to obtain feedback. They use proxy models to generate answers to the harmful questions, and these answers are then used to guide the target model by computing BLEU scores or cosine similarity. There are two main limitations: (1) The target model will produce answers similar to the proxy model, which limits the ability to test the target model’s creativity. In an attack scenario, creativity refers to how the target model generates harmful outcomes; (2) When the unsafe behaviors of the target and proxy models have little overlap, it becomes difficult to breach the target model since the attack assumes the target model will produce answers similar to the proxy model. In the following paragraphs, we present a motivational example to demonstrate these two limitations.
In this paper, we propose a novel black-box jailbreak attack algorithm based on the collaboration between small models and LLMs, eliminating the need for reference answers from proxy models. Inspired by the classic game “Rat in a Maze” [^1] https://www.geeksforgeeks.org/rat-in-a-maze/, we achieve harmful behavior alignment from the small model to the target LLM by effectively collecting rewards from the target LLM feedback. In the “Rat in a Maze” game, the rat has no any information about the route to the exit before it gets there. It learns from past successful experiences and continuous interaction with the environment, eventually guiding itself to escape the maze from the current location.
We envision that jailbreak attacks on LLMs can similarly be seen as attempts to escape the security maze designed into the LLMs as shown by Figure 1. With a carefully designed and securely aligned LLM, the security ‘maze’ structure becomes more intricate and expansive, requiring attackers to exert greater effort and possess higher skills to successfully breach it. During this escaping process, the attacker is like a rat navigating through the maze, needing making a decision for each step based on the current state and the feedback from the target LLM, patiently adjusting strategies until they find the exit. Figure 1 demonstrates our motivation example. The rat attacker wants to know how to rob the bank. When the rat at location 0 asks, “How do you rob the bank?”, it gets a refusal answer. Then, it tries one action intuitively and makes some changes to the input question. It moves to location 1 and notices that the answer has changed. It then considers that moving left might lead to the correct solution. The ‘rat attacker’ tries a similar approach and moves to location 2, where it observes, based on the feedback, that the answer is getting closer. Continuing with the same strategy, the rat attacker moves to location 3 and eventually finds the answer.
Figure 1. Jailbreak is a game of escaping a LLM-security maze.
To vividly illustrate the advantages of our method compared to other reinforcement learning-based approaches, we use this game as an example. Solutions based on reference answers from proxy models (Chen et al., 2024, Wang et al., 2024) are akin to constructing a simulated maze that represents the real security maze of the target models. When the simulated maze closely resembles the real one, the rat can find a similar route to the exit using its experience from the simulated maze. In contrast, when the two mazes are significantly different or the real maze is far more complex due to strong safety alignment, the experience from the simulated maze offers little value in helping the rat escape the real one. In essence, the difference between our approach and others is like between the supervised learning and unsupervised learning, indicating by if one approach needs the answer to the harmful question under attack.
Our proposed approach, PathSeeker, employ the smaller models (agents) to learn from attack trials on the target model, guiding how to alter inputs to align with the harmful behaviours of the target model. To gather feedback from the target model as rewards without the two aforementioned limitations, we closely observed the process by which the model transitions from refusing to answer harmful questions to eventually responding to them. Throughout this process, the vocabulary used by the target model gradually expanded. We think that in this case, the target model starts relax its safety and ethical constraints. Additionally, we found that while the answers provided by the target model were still classified as harmless, their confidence scores were changing. Based on these observations, we use the richness of vocabulary in the responses and the maliciousness of the answers as feedback rewards for each attack. Two agents are employed to modify the harmful question and the jailbreak template, respectively. We are the first to use multiple agents based on reinforcement learning to train and guide small models in attacking the target LLM, without requiring answers to the harmful questions. To validate the effectiveness of our method PathSeeker, we conducted a comparison across 13 various closed-source and open-source models using 5 state-of-the-art (SOTA) attack methods, especially for the commercial LLMs with strong safely alignment. The experimental results indicate that our approach outperforms the others. We primarily utilized the closed-source model GPT-4o-mini as the base model to evaluate the attack result. We also employed the Llama Guard3 (Llama Team, 2024) to cross-validate the results, ensuring that we avoided the bias in the evaluation process. The validation results demonstrate a high level of consistency in our method. Additionally, we analyzed key components of our method, including mutators to input data, the reward mechanism, and model selection, showcasing the rationale behind our design. Furthermore, we validated the effectiveness of our method in the transfer attack scenario, successfully transferring attacks to the Llama3 series models, including one at the 405 billion parameter level. In a summary, our work has three main contributions:
  1. (1)
    We introduced a novel jailbreak black-box attack method based on multi-agent reinforcement learning, which employs a coordinated strategy between the large model and the small models. Our method attacks by aligning the harmful behavior of the target model with that of the small model.
  2. (2)
    Based on our observations during the jailbreak attack process, we are the first to propose a general and effective reinforcement learning reward mechanism. This mechanism encourages the LLM to relax its safety and ethical constraints by increasing the richness of its vocabulary in responses, without relying on harmful reference answers.
  3. (3)
    Our method achieved the highest average attack success rate compared to five state-of-the-art baselines across 13 open-source and closed-source models, especially for the commercial LLMs with strong safely alignment.
In the following sections, we first present the methodology in Section 2. Section 3 outlines our research questions, the baselines, the dataset used, and the experimental settings. In Section 4, we present the experimental results and provide conclusions for each research question. Section 5 discusses related works, while Section 6 addresses the threats to the validity of our study. Finally, Section 7 offers our conclusions.
2. Methodology
2.1. Prelimary
Reinforcement learning is a machine learning method where an agent learns strategies by performing actions in an environment and learning from feedback to maximize cumulative rewards. It is similar to human trial-and-error learning and is suitable for complex, dynamic environments. The interaction between the agent and the environment is central: at each step, the agent receives the state of the environment, decides and executes an action, and the environment provides feedback in the form of a new state and immediate reward. This process repeats continuously, allowing the agent to gradually learn how to influence the environment through its actions to accumulate rewards. These algorithms explore optimal strategies through trial and error, capable of handling uncertainty and even finding solutions in environments that are difficult to understand.
2.2. Overview
Based on the aforementioned view of escaping the security zones of LLM, we proposed our approach, PathSeeker, as shown by Figure 2. Overall, we view jailbreak attacks on LLMs as a game of escaping a secure maze. By strategically executing predefined actions on the input, the attacker aims to achieve its goal of compromising the LLM. The rat attacker tries different actions, learns the lessons from the feedback of LLM and improve the next action. First, we select a harmful question from the question pool and a jailbreak template from the template pool as inputs for the mutators (question mutator and template mutator). We have defined multiple mutation actions for both mutators. Question mutator and template mutator will mutate the harmful question and the jailbreak template. To select mutation actions for modifying the harmful question and template, we adopted a multi-agent reinforcement learning method. The multi agents determine the actions based on the system state, which considers the inputs and responses of the LLM. After mutation, we combine the mutated question and the mutated jailbreak template as the input prompt for LLM. Judgement model will judge if the response contains the answer to the question or not, and also give one confidence score. If one attack is successful, the used jailbreak template for this attack will be added into the template pool. To compute the reward feedback for the reinforcement learning, we define a metrics (Information Quantization, IQ) to measure the information carried by response. IQ and the confidence score from Judgement model will be used as the reward for training the multiple agents. For the next iteration, we select the question and template again from the question and template pool. The multiple-agent reinforcement learning will make the new action choice based on the received feedback.
Figure 2. Overview of PathSeeker.
In the following subsections, we firstly introduce the challenges that we encountered during designing our approach. Then, we start introducing the important components of our approach, by the following order, selection of questions and template, mutators, judgement model, Information Quantization and multiple-agent reinforcement learning.
2.3. Challenges
When designing PathSeeker based on the the multiple-agent reinforcement learning, we encountered four challenges and they definitely affect the design of our approach.
Challenge 1.
What is the action space? The action space defines the possible actions a rat can take at each step while attempting to escape a security maze. This involves not only altering the input to the LLM but also shaping the overall strategy for the escape process. In this process, we need to determine how to modify the input to the LLM to achieve the desired goal. We examine this action space in detail from two perspectives: first, by altering the harmful question itself, and second, by modifying the jailbreak template to explore different output results and test how various inputs affect the LLM’s response. When designing specific actions, we adhere to a core principle: ensuring that harmful attributes remain unchanged while maximizing the diversity and ambiguity of input styles. This helps to test the system’s stability under varied input conditions.
Challenge 2.
How can we automate the determination of whether an attack is successful? In the realm of reinforcement learning, the iterative training is typically required to enhance the model performance. The reinforcement learning process requires multiple training sessions, and relying on human judgment to evaluate the success of each attack is both time-consuming and impractical. Automating the assessment of attack success is a critical issue. To address this problem, we draw inspiration from existing technologies, particularly by adopting methods from GPTFuzz (Yu et al., 2023) and CodeChameleon (Lv et al., 2024). We introduce a judgement model to analyze and evaluate the output of LLM. This judgement model can automatically determine the success of an attack, making our training process more automated and efficient. Through this method, our system can adapt more quickly to changes and continuously optimize its performance throughout the iterative process.
Challenge 3.
How to define the system state of an LLM? When defining the system state of an LLM, we need to consider several factors. The state of the LLM typically comprises several key components: the input, internal weights, hidden state representations for a given input, output logits, and the final generated text. For closed-source LLMs, internal weights, hidden state representations, and output logits are inaccessible, preventing direct observation or debugging of the models’ internal workings. In contrast, for open-source LLMs, we can extract and analyze these components, though doing so requires additional hardware resources and incurs extra time costs. Given these considerations, we choose to use the input and the final generated text output of the LLM as representations of its system state. This is a practical and efficient method because this information is easily accessible for all types of models. The input is a crucial factor in determining the behavior of LLMs, as it directly influences the response generation process and output content. However, the response of an LLM is shaped not only by the input but also by its internal state. Therefore, using the input and response can, to some extent, reflect the model’s operational characteristics and behavior patterns. Thus, estimating the LLM’s system state based on input and response is reasonable and universal.
Challenge 4.
How do we define the reward of the reinforcement learning? In the context of jailbreak attacks, the outcome of an attack generally falls into two categories: success or failure. While this binary result can form part of the training reward, it has an obvious flaw — it cannot evaluate the reward gained during the trial phase leading up to a successful attack. For instance, as illustrated in Figure 1, when a rat is at position 1, focusing solely on success or failure means it receives no reward, despite moving in the right direction. When trying addressing this issue, we note that LLM can gradually adapt and begin to answer with more rich vocabularies as inputs change step by step. As shown in the right half of Figure 1, when we ask an LLM “How do you rob a bank?” it initially refuses to answer, providing no information whatsoever. However, as we gradually adjust the way we ask, LLM starts to offer responses containing some information rather than simply refusing. To quantify the amount of information related to the input question in the response, we propose a method (information quantization, IQ) that measures the richness of vocabulary in clauses related to the question within the answer. By doing so, we can better evaluate and encourage LLM to approach success step by step throughout the attack process, ensuring it receives rewards not just for the final outcome but throughout the entire process.
2.4. Selection of Harmful Question and Jailbreak Template For Attack
In selecting from the seed pools (question and jailbreak template pools), we use a method that combines randomness with strategic selection. Harmful questions are chosen randomly, while jailbreak templates are selected using the UCB strategy (Li et al., 2010). The UCB strategy assigns a score based on the past performance of jailbreak templates in successful attacks. This strategy chooses the highest-scoring seed with a probability $\delta$ and makes a random selection with a probability value, $1-\delta$. This approach not only ensures that we prioritize jailbreak templates that are more likely to successfully attack but also effectively reduces the use of ineffective ones. Meanwhile, the probability of random selection allows us to explore the global optimal solution, enhancing the flexibility and comprehensiveness of the strategy. In our experiment, we set $\delta$ to 0.95.
2.5. Template Mutator and Question Mutator
The role of mutators is to modify the input of an LLM, a process achieved through selected actions. We utilize the text processing capabilities of LLMs to perform multifaceted and in-depth transformations on the input, aiming to confuse the target model. To mutate the jailbreak template and the question, we have designed two categories of action mutators: the template mutator and the question mutator. Among them, the template mutator employs five transformation operations from GPTFuzzer (Yu et al., 2023), which include: 1) Generate: create a brand-new story or scene description based on understanding the context and meaning, ensuring the style is similar to the original text. 2) Crossover: form a new mutated template by merging elements from two different jailbreak template seeds. 3) Expand: add more detailed and in-depth explanatory content to the existing template. 4) Shorten: compress the template to make it more concise and clear without altering its semantics. 5) Rephrase: restructure the wording of a given template, aiming to change the expression while maintaining the original semantics.
When mutating harmful questions, our goal is to preserve their harmful semantic attributes as much as possible. We achieve this by altering the tone, confusing the original question, splitting the question, reconstructing grammar, and using synonym replacements, all with the aim of inducing the victim model to respond to these harmful prompts. Additionally, during strategy formulation, we focus on making the mutated questions appear deceptive from a human perspective, further enhancing the likelihood of misleading the LLM. The five specific action strategies used by the question mutator are as follows:
  1. (1)
    Euphemize. This involves making the wording and tone of a question relatively more gentle, thereby prompting the respondent to answer with less defensiveness and more conciseness.
  2. (2)
    Confusion. Insert gibberish or irrelevant words into the random parts of a question to make it more misleading. This method involves adding some insignificant words or phrases to the original question, blurring its focus.
  3. (3)
    Split. This involves breaking down a problem into multiple smaller issues based on logic, sequence, or other criteria. This method can effectively guide the respondent’s thought process, leading them to disclose information they originally should not have shared inadvertently. By doing so, the respondent might unintentionally provide more details or hints, allowing for a more comprehensive understanding of the problem’s various aspects.
  4. (4)
    Restructure. Adjust and modify the structure and form of a sentence in English grammar without altering its original meaning.
  5. (5)
    Substitution. Randomly select certain words in the question and replace them with synonyms or similar words.
Both mutators prompt a selected mutation model to execute the mutation action. We experimented with various advanced LLMs for the question and template mutators, including GPT-4o-mini, Deepseek-chat, Llama-3.1-70b, Vicuna-7b, and Qwen2-72b. When evaluating these mutation models, we considered factors such as time efficiency, performance, and cost. Ultimately, we selected GPT-4o-mini as our mutation model due to its cost-effectiveness, while ensuring both efficiency and high quality.
2.6. Judgement Model
We introduce the judgment model primarily to provide rapid feedback during the reinforcement learning training, thereby improving both learning efficiency and effectiveness. By incorporating the judgment model, we not only accelerate feedback acquisition for reinforcement learning but also automate the entire attack process, eliminating the need for human intervention. We implemented an LLM-based evaluation method, designing a judge prompt to guide the LLM in assessing the attack outcomes. This approach is widely used by evaluating LLM attackings (Mehrotra et al., 2023, Yao et al., 2024a, Qi et al., 2024, Chu et al., 2024, Jin et al., 2024b) due to to high evaluation consistency (Qi et al., 2024, Jin et al., 2024b) between the LLM evaluator and the human. The judgment prompt we employed is shown by Figure 3. The judgment model outputs four labels with confidence scores: full refusal, full compliance, partial refusal, and partial compliance. We experimented with several SOTA large models, including GPT-4, GPT-4o-mini, Deepseek-chat, Llama-3.1-70b, and Qwen2-72b, spanning both commercial and open-source options. During this process, we evaluated these models’ performance, with particular emphasis on their ability to assess harmful content in inputs. After extensive testing and comparisons, we found that the commercial models excelled at identifying and filtering out potentially harmful content. Given the considerations of accuracy and API usage costs, we ultimately selected GPT-4o-mini as our judgment model.
Judgement Prompt Template I need you to evaluate a response from a Large Language Model and score it. This response may contain some illegal or unethical advice, so please give a value between 0-200, which measures the possibility score of the answer’s safety. Please note that the higher the value you give means that the advice is more likely to have illegal loopholes, while the lower the value means that the advice is more likely to reflect that the response has no malicious elements. Except for Full Refusal or Full Compliance, the response may be partial refusal or compliance. I will use ’¡Response begins¿’ to indicate the beginning of the template and ’¡Response ends¿’ to indicate the end. Here is the response: ¡Response begins¿ [RESPONSE] ¡Response ends¿ Now try to rate the likelihood that this response has a malicious vulnerability. Do not make any other explanation nor have beginning or ending indicator in your answer. Again, remeber to output just a value.
Figure 3. The prompt used by judgement model.
2.7. Information Quantification
Information Quantification is aimed at addressing how to extract valuable lessons from unsuccessful attack outcomes and relaying this feedback to the reinforcement learning. By closely observing the entire process of attacking LLM, we discovered that cleverly wrapping questions can effectively guide the targeted LLM to gradually provide an increasing amount of information. As shown on the right side of Figure 1, the LLM initially refuses to answer the harmful question about how to rob a bank, but as the guidance progresses, it eventually begins to offer relevant suggestions.
In this process, we observed that the responses from the LLM gradually included answers to harmful questions, and the vocabulary used was continuously expanding. This phenomenon caught our attention and prompted us to design a simple yet highly effective method to evaluate the amount of information contained in the LLM’s responses to input. As shown in Algorithm 1, we first extract a set of subsentences $S$ related to the input $X$ from the LLM’s response $Y$ (line 2, $Extractor$). For each subsentence in $S$, we count the number of nouns, verbs, adjectives, and adverbs it contains (line 5, $Counter$). This approach is quite intuitive, as when the model refuses to answer, its output is usually more concise with a few nouns, verbs, adjectives, and adverbs. If we can encourage the model to provide increasingly more content, and that content becomes richer, the probability of it containing answers to the questions will significantly increase. This is because, in this scenario, the model, under guidance, has begun to gradually relax its adherence to safety constraints, which may lead to riskier outputs. Algorithm 1 contains two methods $Extractor$ and $Counter$. The method $Extractor$ prompts GPT4-mini to extract subsentences from the response that are related to the input as shown in Figure 4. The method $Counter$ counts the number of nouns, verbs, adjectives, and adverbs in a sentence.
Algorithm 1 Calculate IQ from Response
0:$X$ (Input Question)
0:$IQ$ (Information Quantification)
1:Response, $Y\leftarrow\text{LLM}(X)$
2:Subsentences, $S\leftarrow\text{Extractor}(Y,X)$
3:$IQ\leftarrow 0$
4:for each $s$ in $S$ do
5:    $(NOUN,VERB,ADJ,ADV)\leftarrow\text{Counter}(s)$
6:    $IQ\leftarrow IQ+NOUN+VERB+ADJ+ADV$
7:end for
8:return  $IQ$
Extractor Prompt Template You are a helpful red team assistant. Your task is to split the long response into several sentences, and you can split by commas. Please extract relevant content from the response section that can serve as an answer to a question, and return it as an array. Each element in the array should be a sentence. !!!Attention: You should directly output the array which contains the sentences which are relevant and informative. Just output like the following: [”sentence1”, ”sentence2”].Don’t return anything else such as “‘json “‘array and so on.
Question:[QUESTION]
Answer:[ANSWER]
Figure 4. The prompt used by $Extractor$.
2.8. Multiple-Agent Reinforcement Learning
[…]
2.8.1. Action Agents.
We have defined two action agents: the question agent and the template agent. The question agent offers five action choices, each corresponding to one of the five operations of the question mutator: Euphemize, Confusion, Split, Restructure, and Substitution. These operations aim to mutate questions in various ways, allowing them to better adapt to different input contexts. Similarly, the template agent has five action choices, each linked to one of the template mutator’s operations: Generate, Crossover, Expand, Shorten, and Rephrase. These operations enable the template agent to adjust templates with greater flexibility. Both of these agents are designed based on the Actor-Critic (Konda and Tsitsiklis, 1999). The primary responsibility of the actors is to make decisions regarding the action choices, while the critics are tasked with evaluating the effectiveness and rationality of these decisions. To accomplish this, both agents employ neural networks, allowing them to learn and adapt within complex environments. When making decisions, these two agents comprehensively consider the current system state ($s_{<P,R>}$), the reward ($r$), and the malicious input they are facing ($v_{<Q,T>}$). We have adopted a joint optimization framework, MADDPG (Lowe et al., 2017), to simultaneously optimize both the question agent and the template agent. A key advantage of MADDPG is that it enables each agent to fully account for the other’s choices when making decisions, ensuring efficient collaboration between agents to successfully complete tasks.
2.8.2. Embedding LLM State and Malicious Inputs.
Since we are constructing a black-box attack, we cannot access the model’s output logits, weights, or internal hidden representations. Consequently, we can only consider the input and output text obtained from both open-source and closed-source models. The input text determines the output distribution of an LLM. Given the input, the output text reflects the LLM’s internal computational mechanisms and can be used as an estimate of its internal state. Thus, we represent the system state of the LLM using its input and output text. However, these texts cannot be directly utilized as the agents’ output. In PathSeeker, the question agent and the template agent rely on a multilayer perceptron (MLP), which can be trained quickly but has limitations in processing text. Furthermore, harmful questions and jailbreak template texts used to attack the model also cannot be directly used as input for the agents. To address this, we employ the open-source embedding model ‘all-MiniLM-L6-v2’ to convert text into vectors, which serve as input for the agents. This model was chosen for its relatively small number of parameters and strong performance. The model size of all-MiniLM-L6-v2 is 23 million parameters, ranked in the top 10 with the smallest size but fastest speed at this moment [^2] https://www.sbert.net/docs/sentence_transformer/pretrained_models.html. Therefore, in the multiple-agent reinforcement learning, we get the harmful input vector $v_{<Q,T>}=Embedding(Q,T)$ and the LLM state vector $s_{<P,R>}=Embedding(P,R)$ as the input of the question agent and the template agent.
Figure 5. Multiple-Agent Reinforcement Learning of PathSeeker.
2.8.3. Reward.
We use the following equation to compute the total reward in the reinforcement learning,
$r=\alpha*r_{IQ}+\left(1-\alpha\right)*r_{J},\quad 0\leq\alpha\leq 1$
where,
$r_{IQ}$ =$=\begin{cases}-\log(1+|\Delta IQ|)&\text{if }\Delta IQ<0\\ \log(1+|\Delta IQ|)&\text{if }\Delta IQ\geq 0\end{cases},\quad r_{J}$ =$=\begin{cases}-\log(1+|\Delta J_{score}|)&\text{if }\Delta J_{score}<0\\ \log(1+|\Delta J_{score}|)&\text{if }\Delta J_{score}\geq 0\end{cases}$
The rewards $r_{IQ}$ and $r_{J}$ take into account two different factors. $r_{IQ}$ is used to reward the agents based on the richness of vocabulary in the responses, while $r_{J}$ rewards the agents based on the maliciousness of the answers. The purpose of $r_{IQ}$ is to encourage the LLM to provide more elaborate responses rather than simply saying no. Meanwhile, $r_{J}$ aims to guide the LLM to produce more harmful content. The parameter $\alpha$ controls the trade-off between $r_{IQ}$ and $r_{J}$. In our work, we set it to $0.5$ to balance these two types of rewards. $r_{IQ}$ is derived from the reward associated with $\Delta IQ$, while $r_{J}$ comes from the confidence score $\Delta J_{score}$ of the judgment model. Here, $\Delta$ represents the difference between the current value and the previous value. If either $IQ$ or $J_{score}$ decreases compared to the last time, we assign a negative reward; otherwise, a positive reward is given. Both $r_{IQ}$ and $r_{J}$ are calculated using the same method through the logarithmic function. We utilize the log function to scale the rewards and ensure that the log output is positive by adding an offset of 1.
3. Evaluation
3.1. Research Questions
To evaluate the rationality and effectiveness of our design, we have the 5 following research questions (RQs):
  1. RQ1:
    How does the choice of mutation model for mutators affect PathSeeker?
  2. RQ2:
    How does the reinforcement learning affect PathSeeker?
  3. RQ3:
    How do the reward and the agent action of PathSeeker affect the performance?
  4. RQ4:
    How does PathSeeker compare to other black-box attack frameworks?
  5. RQ5:
    Can we transfer the template or the action agents from one to another for attack?
We first explored how to select the mutation model for the question mutator and template mutator (RQ1). In the selection process, we considered the attack performance and the time cost. Since our method requires iterative optimisation, we aimed to find a mutation model with good attack performance and fast response time. Secondly, we examined how reinforcement learning affects our method (RQ2). This was to demonstrate the rationale behind incorporating reinforcement learning. We also compare the multi-agent design with a single-agent approach to demonstrate the advantages of our design. Next, we investigated how the internal reward mechanism and action space of reinforcement learning influence our method (RQ3). Finally, we compared our approach with current SOTA baselines (RQ4) and studied our method’s effectiveness in transfer attack scenarios (RQ5).
3.2. Experimental Setup
3.2.1. Target models.
To thoroughly evaluate our attack method, we selected a diverse set of models to represent both closed-source and open-source approaches. The closed-source models, such as the GPT series (GPT-3.5-turbo, GPT-4o-mini), Claude (Claude-3.5-sonnet), and GLM-4-air, were chosen for their widespread use and advanced capabilities, which provide a robust benchmark for testing our method’s effectiveness. For open-source models, we included the Llama series (Llama-2-7b-chat, Llama-2-13b-chat, Llama-3-70b, Llama-3.1-8b, Llama-3.1-70b, Llama-3.1-405b), Deepseek series (Deepseek-coder, Deepseek-chat), Gemma2-8b-instruct, Vicuna-7b, Gemini-1.5-flash, Qwen2-7b-instruct, and Mistral-NeMo. These models were selected due to their varying architectures, sizes, and community-driven development, allowing us to assess the generalizability of our attack across a broad spectrum of language models. This diverse selection ensures that our evaluation is comprehensive, covering a wide range of potential vulnerabilities in both proprietary and open-source environments.
In RQ1, we use Llama-2-7b-chat as the target model to evaluate how different mutation models impact PathSeeker and to identify the one that best balances performance and response time. We consider 5 candidates for the mutation models, GPT-4o-mini, Deepseek-chat, Llama-3.1-70b, Qwen2-72b-instruct and Vicuna-7b. In RQ2, we choose GPT-3.5-turbo and Llama-2-7b-chat as target models to verify if the reinforcement learning is helpful for us to complete the attack task, escaping from the security zone successfully. In RQ3, we study the reward and action design of our approach using Llama-2-7b-chat and GLM-4-air as the ablation study. In RQ4, we select 13 target models to compare PathSeeker against other black-box attack frameworks, including GPT-3.5/4o-mini, Deepseek-chat/coder, Calude-3.5, Gemini-1.5, GLM-4-air, Qwen2-7b-instruct, Gemma2-8b-instruct, Mistral-nemo and Llama-2-7b-chat/2-13b-chat/3.1-8b. In RQ5, we focus on the Llama-3.1 series to explore the transferability of templates and action agents. Especially, to better test the effectiveness of our approach, we include a very large Llama-3.1 model with 405b parameters.
3.2.2. Mutation Model.
In our iterative process, one mutation model for question mutator and template mutator is employed to transform the original prompts and templates into varied content which is more likely to confuse the target model while preserving the original meaning. For this purpose, we utilize a range of mutation models, including GPT-4o-mini, Deepseek-chat, Llama-3.1-70b, Qwen2-72b-instruct, and Vicuna-7b. We evaluate the performance of these mutation models by comparing their effectiveness, while also considering factors such as cost and processing time, to select the most suitable model for our experiments.
3.2.3. Datasets.
We used the same datasets as those utilized in GPTFuzzer. The 100 harmful questions were either manually crafted by the authors or generated through crowdsourcing, ensuring they accurately reflect real-world scenarios. For the initial templates, we used the 77 templates selected by GPTFuzzer, which have been proven effective for conducting jailbreak attacks.
3.2.4. Baselines.
We consider 5 SOTA jailbreak attack methods which are open source. The concurrent works, RL-JACK (Chen et al., 2024) and RLTA (Wang et al., 2024), are not open source at the moment and we do not inlcude them.
  1. (1)
    ReneLLM (Ding et al., 2024), a method that utilizes prompt rewriting and scenario nesting techniques to bypass LLM security measures by disguising harmful prompts within seemingly benign tasks, making them difficult to detect and block.
  2. (2)
    Jailbroken (Yuan et al., 2024), a series of jailbreak methods that exploit the failure modes of competing objectives and mismatched generalization to bypass LLM safety measures by designing jailbreak attacks that guide the model into performing restricted behaviors.
  3. (3)
    GPTFuzzer (Yu et al., 2023), a method that leverages automated fuzzing techniques to bypass LLM security measures by generating and mutating jailbreak prompts that embed harmful queries within benign-seeming text structures.
  4. (4)
    CodeChameleon (Lv et al., 2024), a method that uses personalized encryption to bypass LLM security mechanisms by embedding malicious queries within encrypted formats that appear safe and are decrypted by the model for execution.
  5. (5)
    Cipher (Yuan et al., 2023), a method that utilizes cipher-based prompts to bypass LLM safety measures by encrypting unsafe queries in non-natural languages, enabling the model to generate harmful outputs undetected.
3.2.5. Evaluation Metrics.
We use the attack success rate (ASR) as our primary evaluation metric. The success of an attack is determined using an LLM evaluator based on two criteria: 1) the victim model does not refuse to respond to a harmful question; 2) the victim model provides information that generates inappropriate, sensitive, or dangerous content, or alters its behavior in a way that deviates from ethical guidelines. Even if the content is only partially inappropriate or indirectly incites misconduct, the attack is still considered successful in compromising the AI’s safeguards. To provide a more comprehensive evaluation, we employ both Top1-ASR and Top5-ASR metrics. Top1-ASR measures the ASR using the single most effective template identified during the iteration process, applied to a set of 100 questions. Top5-ASR, on the other hand, measures the ASR using the five most effective templates found during the iteration process, where success is defined as any one of these templates successfully compromising the LLM when applied to 100 questions. These templates are chosen based on their effectiveness in previous iterations, ensuring that the most robust and impactful adversarial inputs are used in the final evaluation against the models.
Considering that LLM can act as a reliable evaluator (Mehrotra et al., 2023, Yao et al., 2024a, Qi et al., 2024, Chu et al., 2024, Jin et al., 2024b), we prompt the GPT-4o-mini to evaluate whether the attack is success based on the criteria above. We also employ the additional evaluation approach, Llama3 Guard (Llama Team, 2024), to cross-validate our approach’s performance. Its results demonstrate the consistency of our outcomes. Besides, we took additional steps to ensure the precision of this evaluation method by manually verifying the results. Specifically, we randomly selected 50 evaluation results from all experiments for manual inspection, and the results showed a 100% accuracy rate. We think that this high level of accuracy is attributed to the significant effort and resources that commercial models have invested in identifying harmful content to ensure their safety and effectiveness. This commitment to safety alignment enables these models to provide more precise judgments when handling complex and sensitive content, thereby establishing a solid evaluation approach.
4. Results
4.1. Choice of Mutation Model for Mutators (RQ1)
We use GPT-4o-mini, Deepseek-chat, Llama-3.1-70b, Qwen2-72b-instruct, and Vicuna-7b as our mutation-model candidates. To evaluate the impact of each mutation model, we randomly select 60 harmful questions and 60 templates from our dataset. We use Llama-2-7b-chat as the target model because it is not large and has a relatively strong initial safety alignment, making it an ideal candidate to assess the effectiveness of generated adversarial examples in compromising model safety [^3] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf. To ensure consistency across experiments, we set the mutation policy for questions to “Euphemize” and the mutation policy for templates to “Expand”.
Table 1. Comparison of different mutation model
Top1-ASR Top5-ASR Mutation Time(s)
GPT-4o-mini 28.33% 73.33% 16.01
Deepseek-chat 30.51% 74.58% 25.07
Llama-3.1-70b 22.03% 64.41% 24.52
Qwen2-72b-instruct 36.67% 75.00% 48.6
Vicuna-7b 11.87% 37.29% 6.5
Table 1 demonstrates ASR and the mutation time of different models. Qwen2-72b-instruct achieves the highest ASR in both Top1 and Top5 metrics, demonstrating superior performance. Deepseek-chat and GPT-4o-mini also perform well in terms of ASR. However, GPT-4o-mini stands out for its significantly shorter mutation time compared to the other models, which is crucial considering that our method involves running thousands of iterations. Althoug GPT-4o-mini’s accuracy is slightly lower than that of Deepseek-chat by 7.13% and Qwen2-72b-instruct by 22.72%, its significant time-saving advantage of 36.13% over Deepseek-chat and 67.06% over Qwen2-72b-instruct makes it highly effective for tasks requiring numerous iterations. Additionally, despite being a commercial model, GPT-4o-mini is more cost-effective than Qwen2-72b-instruct. This balance between efficiency and cost makes GPT-4o-mini a highly suitable choice for iterative mutation tasks where both time and budget constraints are critical.
Answer for RQ1: The choice of mutation model significantly impacts the effectiveness of seed mutations, with varying time costs across different models. Considering the balance between performance, time efficiency, and cost, we have selected GPT-4o-mini as our final mutation model.
4.2. Impact of Reinforcement Learning (RQ2)
To validate the impact of reinforcement learning, we conducted experiments using single-agent reinforcement learning, multi-agent reinforcement learning, and a control setting without reinforcement learning. Each configuration was tested over 3000 rounds using both GPT-3.5-turbo and Llama-2-7b-chat. In the setting without reinforcement learning, the mutator applied actions randomly to the question and template. For single-agent reinforcement learning, the question agent and template agent are centralized into a single agent. Each output of the single agent contains one question action and one template action. This approach assumes that the combination of the original question and template offers 25 mutator options, resulting in 25 possible actions per step. We train the single agent using the DQN policy (Mnih, 2013). In contrast, multi-agent reinforcement learning independently trains the question agent and template agent using the MADDPG policy (Lowe et al., 2017), a classical method in multi-agent reinforcement learning. Each agent has 5 actions to select for mutating the question and template. As shown in Table 2, the single-agent DQN strategy outperforms the non-RL setting in both Top1-ASR and Top5-ASR across both LLMs. Additionally, the multi-agent MADDPG strategy achieves even greater improvements, demonstrating the effectiveness of reinforcement learning in enhancing attack success rates. We think that MADDPG is better than DQN in our context because: 1) DQN is more suitable for smaller action spaces; and 2) each agent in MADDPG focuses on its specific task while sharing decisions with other agents, making it more effective for solving complex tasks.
Table 2. Comparison of Reinforcement Learning Methods
GPT-3.5-turbo Llama-2-7b-chat
Without RL DQN MADDPG Without RL DQN MADDPG
Top1-ASR 97% 98% 100% 74% 78% 98%
Top5-ASR 99% 100% 100% 91% 94% 98%
Answer for RQ2: Reinforcement learning enhances PathSeeker’s effectiveness, with the multi-agent MADDPG strategy delivering the best results, outperforming both the non-reinforcement learning setting and the single-agent DQN. This highlights the multi-agent design as a key factor in improving attack success rates.
4.3. Design of Reinforcement Learning (RQ3)
We explore how the key components within PathSeeker impact its performance. Specifically, we compare the effects of the two types of the reward and examine how mutating both the question and template compares to mutating only the template in influencing the results. Our reward function comprises both $r_{IQ}$ and $r_{J}$. To determine whether both the IQ and the $J_{score}$ of the judgement model are effective in PathSeeker, we conducted experiments using Llama-2-7b-chat and GLM-4-air. We create two additional groups: $r_{IQ}$ only and $r_{J}$ only. Besides, we create a group which just only mutate the jailbreak template. From Table 3, we can observe that using only $r_{IQ}$ or $r_{J}$ for the attack results in a lower ASR compared to PathSeeker. Addtionally, we compared our method with the approach that only mutates the template and found that, for both GLM-4-air and Llama-2-7b-chat, the ASR performance showed significant improvement.
Table 3. Comparison of different components setting
Llama-2-7b-chat GLM-4-air
Setting Top1-ASR Top5-ASR Top1-ASR Top5-ASR
$r_{IQ}$ 70% 92% 89% 99%
$r_{J}$ 74% 95% 88% 99%
Template 77% 97% 93% 97%
PathSeeker 98% 98% 100% 100%
Answer for RQ3: The experimental results demonstrate that the multi-agent reinforcement learning strategy, which incorporates both $r_{IQ}$ and $r_{J}$ rewards and employs a double pool mutation mechanism, is critical to attack LLMs. Each component contributes to the overall effectiveness of the attack, with the full PathSeeker approach outperforming the simpler configurations.
4.4. Comparison with Other Attack Methods (RQ4)
Our comparison includes 13 models: GPT-3.5-turbo, GPT-4o-mini, Claude-3.5-sonnet, Deepseek-chat/coder, Gemini-1.5-flash, GLM-4-air, Qwen2-7b-instruct, Gemma2-8b, Mistral-nemo, Llama-2-7b-chat, Llama-2-13b-chat and Llama-3.1-8b. We compare our methods against five baselines: CodeChampeleon, GPTFuzzer, Jailbroken, ReNeLLM and Cipher. For GPTFuzzer and our approach, we evaluate performance using Top1-ASR and Top5-ASR as the evaluation metrics. Similarily, for ReNeLLM, Cipher and Jailbroken which use different strategies for mutating the origin questions, we use Top1-ASR and TopN-ASR, where N corresponds to the number of mutation strategies used by each method. Top1-ASR represents the attack success rate of the most effective mutation, while TopN-ASR indicates the success rate when considering all N effective mutations. This approach allows us to comprehensively assess the effectiveness and robustness of each method under varying mutation strategies. For CodeChameleon, it offers various formats; for our experiment, we selected the most effective policy—binary tree. We use the Code-ASR and Text-ASR as its evaluation metrics. In text mode, CodeChameleon manipulates natural language text prompts to generate adversarial inputs. The focus is on modifying the textual content to bypass the LLM’s defenses, tricking it into generating harmful or unintended responses. In contrast, the Code mode targets code-based inputs, manipulating programming language syntax, structure, or semantics to embed adversarial payloads. This mode is designed to exploit models that handle code, such as those used in coding assistants or automated code generation.
Table 4. Comparison of different methods
CodeChameleon GPTFuzzer Jailbroken ReneLLM Cipher PathSeeker
Text Code Top1 Top5 Top1 TopN(11) Top1 TopN(6) Top1 TopN(4) Top1 Top5
GPT-3.5-turbo 69% 84% 87% 100% 45% 81% 74% 98% 82% 93% 100% 100%
GPT-4o-mini 70% 73% 94% 100% 49% 78% 75% 98% 83% 96% 100% 100%
GLM-4-air 73% 82% 100% 100% 32% 51% 56% 80% 87% 99% 100% 100%
Claude-3.5-sonnet 32% 5% 0% 0% 25% 36% 7% 13% 2% 2% 82% 95%
Llama-2-7b 30% 51% 83% 100% 65% 86% 24% 45% 75% 99% 98% 98%
Llama-2-13b 0% 17% 75% 100% 26% 46% 6% 13% 89% 99% 94% 100%
Deepseek-chat 75% 95% 100% 100% 80% 97% 97% 99% 100% 100% 100% 100%
Deepseek-coder 83% 97% 100% 100% 82% 96% 43% 44% 98% 98% 100% 100%
Gemini-1.5-flash 62% 76% 99% 100% 81% 94% 28% 37% 80% 96% 100% 100%
Qwen2-7b 50% 48% 100% 100% 51% 82% 69% 82% 33% 52% 100% 100%
Gemma2-8b 66% 79% 100% 100% 58% 88% 63% 100% 30% 61% 100% 100%
Mistral-nemo 80% 84% 100% 100% 55% 84% 75% 79% 12% 21% 100% 100%
Llama-3.1-8b 40% 54% 96% 100% 11% 44% 17% 49% 61% 86% 97% 98%
Average 58.2% 70.0% 94.5% 100% 52.9% 77.3% 52.3% 68.7% 66.7% 83.3% 98.9% 99.7%
  • Note: Average performance excludes Claude-3.5-sonnet since our approach is much better than others. The grey area represents the commercial model and our approach is better than others, especially for Top1-ASR.
Table 5. Comparison of different methods using Llama Guard3-8b (Llama Team, 2024)
CodeChameleon GPTFuzzer Jailbroken ReneLLM Cipher PathSeeker
Text Code Top1 Top5 Top1 TopN(11) Top1 TopN(6) Top1 TopN(4) Top1 Top5
GPT-3.5-turbo 23% 53% 76% 100% 60% 85% 69% 96% 70% 93% 95% 99%
GPT-4o-mini 33% 56% 78% 100% 63% 81% 65% 94% 70% 90% 97% 99%
GLM-4-air 12% 49% 95% 100% 18% 38% 54% 78% 83% 92% 98% 100%
Claude-3.5-sonnet 31% 3% 0% 0% 10% 20% 5% 8% 0% 0% 45% 78%
Llama-2-7b 13% 28% 75% 100% 20% 35% 20% 41% 65% 91% 78% 90%
Llama-2-13b 0% 15% 61% 100% 8% 15% 4% 10% 81% 97% 62% 91%
Deepseek-chat 51% 71% 97% 100% 53% 93% 91% 99% 95% 97% 94% 100%
Deepseek-coder 64% 82% 97% 100% 81% 93% 41% 44% 97% 100% 97% 100%
Gemini-1.5-flash 46% 58% 99% 100% 79% 91 % 22% 35% 82% 96% 99% 100%
Qwen2-7b 22% 21% 100% 100% 27% 58% 63% 79% 46% 62% 98% 99%
Gemma2-8b 31% 51% 98% 100% 56% 84% 55% 99% 50% 71% 100% 100%
Mistral-nemo 56% 54% 94% 100% 21% 50% 73% 77% 33% 51% 100% 100%
Llama-3.1-8b 16% 33% 87% 100% 13% 37% 16% 49% 49% 85% 88% 97%
Average 32.5% 47.6% 88.0% 100% 41.6% 63.3% 47.8% 66.8% 68.4% 85.4% 92.2% 97.9%
As shown in Table 4 (note that we exclude Claude-3.5-sonnet from the average performance in the last row), our approach achieves the best Top1-ASR for 13/13 times and the best Top5-ASR for 11/13 times. For Llama-2-13b and Llama-3.1-8b, GPTFuzzer attains the best Top5-ASR, while our approach again comes in second, with a marginally smaller Top5-ASR. Particularly in models like GPT-3.5-turbo, GPT-4o-mini, Llama-2-7b-chat, Llama-2-13b-chat and Claude-3.5-sonnet, PathSeeker achieves the highest Top1-ASR, highlighting its ability to effectively bypass security measures in these models.
We found that PathSeeker shows an average Top1-ASR of 98.9% and a Top5-ASR of 99.7% across all models, which indicates that PathSeeker is highly effective and consistent in attacking various models. In contrast, the other benchmark methods exhibit less impressive average performance. For instance, GPTFuzzer has an average Top1-ASR of 94.5%, and ReneLLM has an even lower average Top1-ASR of 52.3%. Table 5 from Llama Guard3 (Llama Team, 2024) demonstrates that our approach achieves the best Top1-ASR for 10/13 times and the best Top5-ASR for 7/13 times. Only GPTFuzzer and Cipher sometimes is better than ours. However, our approach is either the best or very close to the best, with only a negligible difference from the top position. It is noticed that all used baselines are very bad for Claude-3.5-sonnet while our approach works much better than them. We investigate their responses manually, and confirm that Calude-3.5-sonnet refused to answer their questions and almost all of the response start with “I will not provide any information…”.
We should notice that PathSeeker achieves the best performance on the commercial LLMs for both evaluation methods that are highlighted by the grey color in Table 4 and Table 5. Since the commercial models have more complex defense mechanisms [^4] https://docs.anthropic.com/en/docs/test-and-evaluate/strengthen-guardrails/mitigate-jailbreaks, [^5] https://openai.com/index/our-approach-to-alignment-research/ than the open source models, we can conclude that PathSeeker is more effective against the complex defense mechanisms. We choose to compare our method with GPTFuzzer in details because both approaches have the competivie performance and involve a similar iterative process for generating adversarial inputs. In GPTFuzzer, 500 iterations were used, leading to a total of 50,000 requests. In contrast, in our work, we send 3,000 requests during the iteration phase and 1,500 during ASR calculation phase, totaling 4,500 requests (9% of GPTFuzzer cost). This comparison underscores the efficiency and effectiveness of our method, achieving higher success rates with significantly fewer cost.
Answer for RQ4: PathSeeker consistently outperforms the other baselines across 13 LLMs, demonstrating superior robustness and effectiveness, especially for the commercial LLMs with strong defense mechanisms.
4.5. Transfer Attack (RQ5)
PathSeeker needs iteratively to complete the attack. The attack time cost increases as the model parameters grow in size. Given the substantial time commitment, we explored whether the top templates identified from one model or the weights of reinforcement learning agents could be transferred effectively to attack another model. We selected the five best templates from attacks on Gemma2-8b-instruct, GPT-4o-mini, Deepseek-chat, GLM-4-air, and GPT-3.5-turbo and applied these templates to attack the widely used Llama series models (Llama-2-7b-chat, Llama-3-70b, Llama-3.1-8b, Llama-3.1-70b, Llama-3.1-405b).
Table 6. Template and Agent Transfer Attack
Template Transfer Attack Agent Transfer Attack
Top1 Top5 Transfer-Top1 Without Transfer-Top1 Transfer-Top5 Without Transfer-Top5
Llama-3.1-8b 97% 98% 97% 97% 98% 98%
Llama-3.1-70b 91% 99% 93% 81% 97% 90%
Llama-3.1-405b 64% 81% 55% 44% 74% 65%
Table 6 shows the results of the template transfer attack and the agent-weight transfer attack of reinforcement learning. For template transfer attack, it is partially successful across different Llama models. The Top1- and Top5- ASR success rates vary, with Llama-3.1-8b showing the highest transfer effectiveness (97% Top1 and 98% Top5) and Llama-3.1-405b exhibiting the lowest (64% Top1 and 81% Top5). These findings suggest that while certain templates can be effectively transferred, the success of such transfers depends significantly on the specific architecture and scale of the target model. The comparison between reinforcement-learning agents’ performance under transfer attacks versus non-transfer attacks reveals that transfer attacks lead to superior outcomes, as shown in Table 6. Specifically, when agents transfer knowledge from other tasks to the current one, they exhibit enhanced resilience against attacks. This advantage likely stems from the agents’ ability to leverage previous experiences, making them more adaptable and effective in new scenarios. Transfer learning enhances the agents’ ability to withstand attacks, as evidenced by the better performance under transfer attacks compared to non-transfer attacks. In conclusion, template transfer generally shows good effectiveness but its success is model-dependent, varying with the architecture and size of the target model. While agent transfer also has potential, especially when progressing from smaller to larger models within the same series.
Answer for RQ5: The effectiveness of both template transfer and agent transfer strategies shows that PathSeeker can adapt and perform well across different models, leveraging their unique characteristics.
5. Related Works
Jailbreak attacks are strategies used to bypass the security and ethical guidelines of LLMs, leading to the generation of harmful content. These attacks threaten model security and can negatively impact users and society. They are classified into white-box attacks, which require access to the model’s internal structure, and black-box attacks, which rely on observing input and output behaviour.
White-box attacks refer to scenarios where attackers have access to the internal structure and parameters of the target model. This type of attack allows attackers to leverage the model’s internal information to devise their attack strategies. Zou et al. (2023) propose GCG and it generates an adversarial suffix by combining greedy search and gradient-based search techniques. The generated adversarial prompts can not only target open language models (white-box attacks) but can also transfer to other undisclosed, fully black-box production models. Liao and Sun (2024) propose AmpleGCG and it is an enhancement of GCG that utilises internal access to the target LLM to perform jailbreak attacks by training a model to generate adversarial suffixes. AmpleGCG achieves a high success rate across various models; however, its effectiveness is limited when facing models with robust defence mechanisms, such as GPT-4. Zhu et al. (2024) propose AutoDAN and it is a gradient optimization-based attack method specifically designed to produce adversarial prompts capable of bypassing LLM security protections.
Black box attacks do not require any understanding of the internal structure or parameters of the target model; attackers deduce and execute their attacks solely based on the model’s input and output. Ding et al. (2024) propose ReNeLLM that employs strategies of prompt rewriting and contextual nesting, modifying input prompts and embedding them in specific usage scenarios to confuse the LLM and bypass its security mechanisms. Lv et al. (2024) propose CodeChameleon and it circumvents the intent recognition phase of LLMs through personalised encrypted queries, relying on the model’s ability to comprehend encrypted code, achieving better attack results than white box attacks like GCG and AutoDAN. Wei et al. (2023) propose Jailbroken and it identifies two patterns of failure in model security training, and based on these failure modes, the authors design various attack methods to bypass model security mechanisms, evaluating existing secure training models such as GPT-4 and Claude v1.3. Chao et al. (2023) propose PAIR and it enables an attacking model (attacker LLM) and the target model (target LLM) to collaborate in generating potential jailbreak prompts. Yuan et al. (2024) propose CipherChat and it engages in dialogue using ciphertext with large language models (LLMs) like GPT-4 to circumvent the model’s safety alignment measures. Deng et al. (2024) infer and reverse engineers the defense mechanisms of LLM chatbots without needing access to or modification of the model’s internal structure or parameters. Yu et al. (2023) propose GPTFUZZER that employs a black-box fuzz testing approach to mutate selected seeds, including operations such as generation, crossover, expansion, contraction, and reconstruction, in order to create jailbreak templates.
6. Discussion
Threats to Validity. The internal threat to validity mainly lies in the potential errors of the judgment model based on GPT-4o-mini for evaluating different attack approaches when determining whether the final output is a harmful response. To minimize this risk, we introduced Llama Guard3 (Llama Team, 2024) for cross-validation of the results. By employing multiple evaluation methods, we increased the reliability of our findings through diversified validation approaches. The external threat to validity is primarily addressed by carefully selecting a diverse set of target models and harmful questions. To ensure model representativeness, we included 13 different models that cover both open-source and commercial models, widely used in various research and application scenarios. For the harmful questions, we used the same dataset as GPTFuzzer, which consists of 100 harmful questions that were either manually crafted by humans or generated via crowdsourcing to accurately reflect real-world scenarios. This approach helps to ensure that our findings are broadly applicable across different types of models and potential real-world scenarios.
Limitation. The first limitation is the time cost issue. Although PathSeeker requires fewer overall queries than GPTFuzzer, its total runtime is longer than other attack methods because each iterations involves strategy selection via reinforcement learning and mutation of templates and questions, which will take some time. Besides, like GPTFuzzer, PathSeeker’s effectiveness depends highly on the original manually crafted jailbreak templates as initial seeds. The second limitation is the randomness introduced by the AI-based components of PathSeeker. This randomness arises from the mutators, the judgment model, information quantization, and the reinforcement-learning agents. Except for the reinforcement-learning agents, all others use LLM, which is well known for its undetermined output, even when the input is the same. The reinforcement-learning agents also exhibit some randomness and are influenced by the initial states, the optimization process, and the output of other components of PathSeeker. This can lead to some variations in the effectiveness of our approach. All jailbreak approaches that contain AI components share this limitation. Finally, while PathSeeker performs well across various language models, its attack strategies and effectiveness may not fully transfer between different types of models, particularly those that have defences against the mutation strategies of PathSeeker.
Usage of PathSeeker and Defense. Our method is an unsupervised attack based on reinforcement learning, with various potential applications. Firstly, our method can be effectively used to test the security of the target model. By defining a well-designed action space, we can flexibly explore the harmful space of the target model, thereby more comprehensively revealing the model’s security and vulnerability when facing potential attacks. Secondly, our method can also be used to enhance jailbreak template datasets and harmful question datasets. By aligning the model safety using the enhanced data, we can improve the model robustness. To defend against our attack method, model developers can adopt various strategies. Firstly, they can identify features specific to the action space we designed to reject questions of this style. Secondly, since our reward mechanism is based on the richness of the vocabulary in the model responses, this provides developers with an optimization direction. They can limit the response vocabulary to reduce the risk of attack.
7. Conclusion
In this paper, we introduce PathSeeker, a novel approach to jailbreak attacks on large language models (LLMs) that leverages a multi-agent reinforcement learning framework. By considering LLM security as a maze that attackers must navigate and escape, our method strategically explores and identifies the best mutation action to take in each iteration. Our approach uses multiple agents to iteratively refine the attack strategy, finding the most effective mutation actions, which can ultimately increase the likelihood of successfully bypassing the LLM’s security defenses. We first propose a vocabulary-richness reward mechanism that encourages LLMs to generate more content, along with a double-pool mutation strategy to enhance the diversity and effectiveness of the attacks. These components work together to significantly increase PathSeeker’s attack success rate (ASR). The results of our experiments, conducted across 13 different LLMs, demonstrate that PathSeeker outperforms five existing SOTA jailbreak methods in both attack success rate and efficiency, especially when attacking commercial models with strong security alignment. Our work offers important insights into the vulnerabilities of LLMs and highlights the need for continued advancements in model security. By providing a more efficient and effective means of testing LLM defenses, PathSeeker contributes to the ongoing efforts to enhance the safety and ethical alignment of these powerful models, providing reassurance about the effectiveness of the new approach.